CHARACTERIZATION OF A 1:1 METHANOL-BENZENE COMPLEX USING MATRIX ISOLATION IN-FRARED SPECTROSCOPY

\underline{JAY C. AMICANGELO}, NATALIE C. ROMANO, AND GEOFFREY R. DEMAY, School of Science, Penn State Erie, Erie, PA 16563.

Matrix isolation infrared spectroscopy was used to characterize a 1:1 complex of methanol (CH_{3}OH) with benzene (C_{6}H_{6}). Co-deposition experiments with CH_{3}OH and C_{6}H_{6} were performed at 17 - 20 K using nitrogen and argon as the matrix gases. New infrared bands attributable to the CH_{3}OH-C_{6}H_{6} complex were observed near the O-H and C-O stretching vibrations of CH_{3}OH and near the hydrogen out-of-plane bending vibration of C_{6}H_{6}. The initial identification of the new infrared bands observed was established by performing a conentration study (1:200 to 1:2000 S:M ratios), by comparing the co-deposition spectra with the spectra of the individual monomers, by matrix annealing experiments, and by performing experiments using isotopically labeled methanol (CD_{3}OD) and benzene (C_{6}D_{6}). Quantum chemical calculations were also performed for the CH_{3}OH-C_{6}H_{6} complex using density functional theory and ab initio methods. Two stable minima were found for the complex: one in which the CH_{3}OH is above the C_{6}H_{6} ring with the hydroxyl hydrogen interacting with the π cloud of the ring (H-π complex) and the other in which the CH_{3}OH is in the plane of the C_{6}H_{6} ring with the hydroxyl oxygen interacting with one of the C-H bonds of the ring (CH-O complex). Comparing the calculated shifts of the vibrational frequencies for both complexes to the observed experimental frequency shifts, it is found that the H-π complex is in best agreement with the experimental shifts in both magnitude and direction. Therefore, it is concluded that the geometry of the CH_{3}OH-C_{6}H_{6} complex observed in the matrix isolation experiments is the H-π complex.